# Black hole based general relativistic limit of $f(R)$ theory of gravity

### v1

@@ 1 Introduction
It was expected earlier that orbits of very short period S-stars in the nuclear star cluster near Sgr A* black hole can be used to measure angular momentum ($a$) and quadrupole moment Q of the black hole. But there are potential complications. Detection of such short period stars is associated with perturbation from other compact objects near the black hole (Sadeghian and Will, 2011). Merritt et al. (2010) showed that testing "no-hair" theorem requires stellar orbits as compact as 0.2mpc ($\sim 41$) au.
@@
A promising gravitational alternative to GR, which addresses late time cosmic acceleration (Capozziello and Fang (2002); Nojiri and Odintsov (2003); Carroll et al. (2004); Sotiriou and Faraoni (2010)), and flat rotation curve of galaxies (Capozziello et al. (2007)) without invoking exotic dark energy and dark matter is the $f(R)$ gravity theory. Here the gravitational field equations are obtained from the Einstein-Hilbert action with Lagrangian of the gravitational field modified as $R\longrightarrow f(R)$, $f$ being a function of the Ricci scalar. It is written as (in the unit of $G=c=1$)
@@
Current constraints on Yukawa like fifth force in the GC environment also come from the analysis of shadow cast by the GC black hole (Vagnozzi et al. (2023)) and measurement of Schwarzschild precession of S2 (Abd El Dayem et al. (2025)). Vagnozzi et al. (2023) constrained interesting extension of GR such as string inspired spacetime, alternative theories of gravity, regular black holes, black hole mimickers, wormhole and naked singularity spacetimes through the EHT shadow of Sgr A*.
@@ 4 Results and discussions
Solar system bound on the PPN parameter $\gamma$ is successfully reproduced by $10^{-16}$ eV scalarons. This scalaron mass is tantalizingly close to the one which is compatible with general relativistic periapsis shift of S2 star, size of the GC black hole shadow and advance of perihelia of planets in the solar system. We conclude with an exciting possibility that general relativistic limit of $f(R)$ gravity theory is scale invariant. The scalaron mass scale presented here is likely to give us a new understanding of gravitational physics in a new regime of spacetime.

### v2

@@ 1 Introduction
It was expected earlier that orbits of very short period S-stars in the nuclear star cluster near Sgr A* black hole can be used to measure angular momentum ($a$) and quadrupole moment Q of the black hole. But there are potential complications. Detection of such short period stars is associated with perturbation from other compact objects near the black hole (Sadeghian and Will (2011)). Merritt et al. (2010) showed that testing "no-hair" theorem requires stellar orbits as compact as 0.2mpc ($\sim 41$) au.
@@
A promising gravitational alternative to GR, which addresses late time cosmic acceleration (Capozziello and Fang (2002); Nojiri and Odintsov (2003); Carroll et al. (2004); Sotiriou and Faraoni (2010) and flat rotation curve of galaxies (Capozziello et al. (2007)) without invoking exotic dark energy and dark matter is the $f(R)$ gravity theory. Nojiri and Odintsov (2011) studied compatibility of traditional $f(R)$ and Horava-Lifshitz $f(R)$ gravity in realizing unified description of inflation and dark energy. Possibility of realizing bouncing cosmology along with unification of inflation and late time accelerated expansion of the universe in modified gravity theories has also been reviewed (Nojiri et al. (2017)). Here the gravitational field equations are obtained from the Einstein-Hilbert action with Lagrangian of the gravitational field modified as $R\longrightarrow f(R)$, $f$ being a function of the Ricci scalar. It is written as (in the unit of $G=c=1$)
@@
Current constraints on Yukawa like fifth force in the GC environment also come from the analysis of shadow cast by the GC black hole (Vagnozzi et al. (2023); Nojiri and Odintsov (2025)) and measurement of Schwarzschild precession of S2 (Abd El Dayem et al. (2025)). The study of long range fifth force using near earth asteroids have been previously conducted (Tsai et al. (2023, 2024)). Vagnozzi et al. (2023); Khodadi et al. (2024) constrained interesting extension of GR such as string inspired spacetime, alternative theories of gravity, regular black holes, black hole mimickers, wormhole and naked singularity spacetimes through the EHT shadow of Sgr A*. The study the M87* black hole with respect to its shape and size has been studied in Bambi et al. (2019), where it was found that the Kerr bound on spin is violated.
@@ 4 Results and discussions
Solar system bound on the PPN parameter $\gamma$ is successfully reproduced by $10^{-16}$ eV scalarons. This scalaron mass is tantalizingly close to the one which is compatible with general relativistic periapsis shift of S2 star, size of the GC black hole shadow and advance of perihelia of planets in the solar system. The scalaron mass scale reported here is also compatible with the mass scale of the Yukawa fifth force mediator and ultralight dark matter ($10^{-21}-10^{-15}$ eV) coming from asteroid astrometric data and $10^{-18}-10^{-17}$ eV coming from analysis of the orbit of the asteroid Bennu (Tsai et al. (2023, 2024)). We conclude with an exciting possibility that general relativistic limit of $f(R)$ gravity theory is scale invariant. The scalaron mass scale presented here is likely to give us a new understanding of gravitational physics in a new regime of spacetime.